OpenStudy (anonymous):

Use a half angle formula to fill in the blanks in the identity below: (cos(7x))^2

OpenStudy (anonymous):

the blank is ___________+__________cos(________x)

OpenStudy (anonymous):

plz help me slove this question thanks much

OpenStudy (nikvist):

use \[\cos{14x}=\cos^2{7x}-\sin^2{7x}\]and\[1=\cos^2{7x}+\sin^2{7x}\]